if sin(theta) +sin(phi)=a and cos(theta) + cos (phi)=b, then find cos2(theta) + cos2(phi)

Grade:11

1 Answers

poojitha
6 years ago
sin (theta)+sin(phi)=a,cos(theta)+cos(phi)=b,sin(theta)+0=a,cos(theta)-1=b,cos2(theta)+cos2(phi)                =cos2(theta)+1=cos^2(theta)-sin^2(theta)+1=(b+1)2-a2+1=b2-a2+2b+2